# A hybrid LLM workflow can help identify user privilege related variables in programs of any size

### v1

@@ Abstract
Many programs involves operations and logic manipulating user privileges, which is essential for the security of an organization. Therefore, one common malicious goal of attackers is to obtain or escalate the privileges [1, 2], causing privilege leakage. To protect the program and the organization against privilege leakage attacks, it is important to eliminate the vulnerabilities which can be exploited to achieve such attacks. Unfortunately, while memory vulnerabilities are less challenging to find, logic vulnerabilities are much more imminent, harmful and difficult to identify. Accordingly, many analysts choose to find user privilege related (UPR) variables first as start points to investigate the code where the UPR variables may be used to see if there exists any vulnerabilities, especially the logic ones. In this paper, we introduce a large language model (LLM) workflow that can assist analysts in identifying such UPR variables, which is considered to be a very time-consuming task. Specifically, our tool will audit all the variables in a program and output a UPR score, which is the degree of relationship (closeness) between the variable and user privileges, for each variable. The proposed approach avoids the drawbacks introduced by directly prompting a LLM to find UPR variables by focusing on leverage the LLM at statement level instead of supplying LLM with very long code snippets. Those variables with high UPR scores are essentially potential UPR variables, which should be manually investigated. Our experiments show that using a typical UPR score threshold (i.e., UPR score $>$0.8), the false positive rate (FPR) is only 13.49%, while UPR variable found is significantly more than that of the heuristic based method.
@@ IV-B Workflow Overview
1⃝ First, from the source code files, we generate the PDG for each function through static analysis. Each nodes in the PDG corresponds to a statement expression. 2⃝ The PDGs are then sliced into variable subgraphs, each of which corresponds to a single variable in the program. 3⃝ Next, based on the variable subgraphs, we collect the code statements, creating a set of code statements that are either control and data dependent on or by the variables. 4⃝ This set of code statements are subsequently rated by an LLM, using an initial prompt that is designed upon the purpose of the target program. The ratings shall reflect the significance of the statements in terms of their impacts toward potential non-control data attacks against the target program. 5⃝ Then, the statement ratings will be used to update the variable subgraphs, serving as node score. 6⃝ Finally, we compute the score of the variable by aggregating the node scores using a aggregation algorithm based on the graph topology. To confirm a UPR variable, subsequent manual review is necessary for those with highest ratings.
@@ VIII Conclusion
In this paper, we introduce a large language model (LLM) workflow that can assist analysts in identifying such UPR variables, which is a very time-consuming task. Our tool will analyze all the variables in a program to output a UPR score for each variable, which is the degree of relationship between the variable and user privileges. When our approach is used in real world, those variables with high UPR scores will then be manually investigated by the analyst. Our experiments show that in a typical setting (i.e., UPR score $>$0.8), the false positive rate (FPR) is only 13.49%, while UPR variable found is significantly more than that of the heuristic based method.

### v2

@@ Abstract
Many programs involves operations and logic manipulating user privileges, which is essential for the security of an organization. Therefore, one common malicious goal of attackers is to obtain or escalate the privileges [1, 2], causing privilege leakage. To protect the program and the organization against privilege leakage attacks, it is important to eliminate the vulnerabilities which can be exploited to achieve such attacks. Unfortunately, while memory vulnerabilities are less challenging to find, logic vulnerabilities are much more imminent, harmful and difficult to identify. Accordingly, many analysts choose to find user privilege related (UPR) variables first as start points to investigate the code where the UPR variables may be used to see if there exists any vulnerabilities, especially the logic ones. In this paper, we introduce a large language model (LLM) workflow that can assist analysts in identifying such UPR variables, which is considered to be a very time-consuming task. Specifically, our tool will audit all the variables in a program and output a UPR score, which is the degree of relationship (closeness) between the variable and user privileges, for each variable. The proposed approach avoids the drawbacks introduced by directly prompting a LLM to find UPR variables by focusing on leverage the LLM at statement level instead of supplying LLM with very long code snippets. Those variables with high UPR scores are essentially potential UPR variables, which should be manually investigated. Our experiments show that using a typical UPR score threshold (i.e., UPR score $>$0.8), the false positive rate (FPR) is only 13.49%, while the number of UPR variables found is significantly more than that of the heuristic based method.
@@ IV-B Workflow Overview
① First, from the source code files, we generate the PDG for each function through static analysis. Each nodes in the PDG corresponds to a statement expression. ② The PDGs are then sliced into variable subgraphs, each of which corresponds to a single variable in the program. ③ Next, based on the variable subgraphs, we collect the code statements, creating a set of code statements that are either control and data dependent on or by the variables. ④ This set of code statements are subsequently rated by an LLM, using an initial prompt that is designed upon the purpose of the target program. The ratings shall reflect the significance of the statements in terms of their impacts toward potential non-control data attacks against the target program. ⑤ Then, the statement ratings will be used to update the variable subgraphs, serving as node score. ⑥ Finally, we compute the score of the variable by aggregating the node scores using a aggregation algorithm based on the graph topology. To confirm a UPR variable, subsequent manual review is necessary for those with highest ratings.
@@ VIII Conclusion
In this paper, we introduce a large language model (LLM) workflow that can assist analysts in identifying such UPR variables, which is a very time-consuming task. Our tool will analyze all the variables in a program to output a UPR score for each variable, which is the degree of relationship between the variable and user privileges. When our approach is used in real world, those variables with high UPR scores will then be manually investigated by the analysts. Our experiments show that in a typical setting (i.e., UPR score $>$0.8), the false positive rate (FPR) is only 13.49%, while the number of UPR variables found is significantly more than that of the heuristic based method.